The unemployment rate in a city is 15%. If 7 people from the city are sampled at random, find the probability that fewer than 2 of them are unemployed.
Carry your intermediate computations to at least four decimal places, and round your answer to at least two decimal places.

Answer by Boreal(15235)

Interested in 1 or 0
P(0)=0.85^7=0.3206
P(1)=7(.15)^1(.85)^6=0.3960; 7 different ways to have 1 unemployed.
Sum of those two is 0.7166